# Photon pair generation via down-conversion in III-V semiconductor microrings: modal dispersion and quasi-phase matching

### v1

@@ IV Conclusion and Future Work
We have derived general formulae for the generation rate of photon pairs, and the generation of squeezed light in the lowest-order squeezing (LOS) regime, from the process of SPDC in ring resonator structures. Since our approach treats the scattering losses quantum mechanically, we can also analyze the statistics of the scattered photons. We have focused on microring resonators made of III-V semiconductors, point-coupled to a waveguide, and our approach can easily be extended to multiple input waveguides, and to microrings made from different materials.
With the integrated photonic structures we have considered, where modal dispersion is significant, the quasi-phase matching associated with the tensor nature of the nonlinearity does not play an essential role in obtaining high efficiencies, although it does affect what structures will lead to the highest efficiencies. From our sample calculations, we obtain optimized structures yielding high generation efficiencies of $3.22\times 10^{-5}$ in the pair-regime, and we have detailed how both the rate at which pairs of generated photons exit the channel without loss, and the heralding efficiencies, depend on the coupling between the channel and the resonator.
@@
In future work we will extend this calculation to treat the higher levels of squeezing that are possible, and include third-order processes such as self- and cross-phase modulation that will affect the dynamics of the squeezing. Additionally, because of the predicted pump depletion, this a promising system to study the entanglement between the pump, signal, and idler fields, which can lead to the deterministic generation of non-Gaussian states [22].
@@ Acknowledgements
Appendix A Lab and ring frames
With $\chi^{(2)}(\bm{r})$ the second-order nonlinear susceptibility of the material at $\bm{r}$, take $\chi^{(2)}_{ijk}{(\bm{r})}$ and $\chi^{(2)}_{i^{\prime}j^{\prime}k^{\prime}}(\bm{r})$ to be the components of the tensor in the ring frame and the lab frame respectively, with $i,j,k\in\{x,y,z\}$ and $i^{\prime},j^{\prime},k^{\prime}\in\{x^{\prime},y^{\prime},z^{\prime}\}$ (see Fig. 1). For a zincblende structure grown in the $\hat{z}$ direction, the only non-zero components of ${\chi^{(2)}}{(\bm{r})}$ for $\bm{r}\in$ ring in the lab frame are:
@@ Appendix A Lab and ring frames
Let $\bm{M}_{\phi}$ be the matrix that transforms the coordinates from the lab frame to the ring frame, by rotation of an angle $\phi$:
@@ Appendix B Degenerate SPDC
In this Appendix we treat degenerate SPDC, where both of the generated photons are within one frequency bin labelled by the centre frequency $\omega_{S}$. For variables that differ from those for non-degenerate SPDC, we will include a “hat” for the degenerate case. The Hamiltonian that describes the degenerate SPDC interaction is
@@ B.2 Pulsed excitation
B.2.1 Pair-regime
In the pair-regime using the same approach from Sec. II.5.1, we write the generated ket (similar to [21]):
@@ B.2.1 Pair-regime
where $|\hat{\beta}|^{2}/2$ is the probability of generating a pair of photons. We write the normalized two-photon ket (i.e. $\langle\hat{\mathrm{II}}\,|\,\mathopen{}\hat{\mathrm{II}}\rangle=1$)
@@
$\begin{split}\frac{|\hat{\beta}_{\lambda\lambda^{\prime}}|^{2}}{2}=&\frac{4\pi^{2}E_{P}}{(v_{P}^{\mathrm{ac}})^{2}\hbar^{3}\omega_{0}}\\ &\times\int\mathrm{d}k_{1}\mathrm{d}k_{2}\left|\hat{\mathsf{K}}_{\lambda\lambda^{\prime}}(k_{1},k_{2},X_{P})\phi\left(X_{P}\right)\right|^{2},\end{split}$ (128)
@@
The pair-regime efficiency is given by
^$\begin{split}\hat{\mathcal{E}}_{\mathrm{pulse;pair-regime}}^{SS}&=\frac{|\hat{\beta}|^{2}}{2|\alpha|^{2}},\end{split}$ (130)
@@
[^$\begin{split}[\hat{\mathcal{E}}_{\mathrm{pulse;pair-regime}}^{SS}]_{\lambda\lambda^{\prime}}&=\frac{|\beta_{\lambda\lambda^{\prime}}|^{2}}{2|\alpha|^{2}},\end{split}$ (131)
@@
$\begin{split}\frac{|\hat{\beta}_{\lambda\lambda^{\prime}}|^{2}}{|\hat{\beta}_{\mu\mu^{\prime}}|^{2}}=\frac{[\hat{\mathcal{E}}_{\mathrm{pulse;pair-regime}}^{SS}]_{\lambda\lambda^{\prime}}}{[\hat{\mathcal{E}}_{\mathrm{pulse;pair-regime}}^{SS}]_{\mu\mu^{\prime}}}=\frac{\eta_{S}^{\lambda}\eta_{S}^{\lambda^{\prime}}}{\eta_{S}^{\mu}\eta_{S}^{\mu^{\prime}}}=\frac{\Gamma_{S}^{\lambda}\Gamma_{S}^{\lambda^{\prime}}}{\Gamma_{S}^{\mu}\Gamma_{S}^{\mu^{\prime}}}.\end{split}$ (132)

### v3

@@ IV Conclusion and Future Work
We have derived general formulas for the generation rate of photon pairs, and the generation of squeezed light in the lowest-order squeezing (LOS) regime, from the process of SPDC in ring resonator structures. Since our approach treats the scattering losses quantum mechanically, we can also analyze the statistics of the scattered photons. We have focused on microring resonators made of III-V semiconductors, point-coupled to a waveguide, and our approach can easily be extended to multiple input waveguides, and to microrings made from different materials.
With the integrated photonic structures we have considered, where modal dispersion is significant, the quasi-phase-matching associated with the tensor nature of the nonlinearity does not play an essential role in obtaining high efficiencies, although it does affect what structures will lead to the highest efficiencies. From our sample calculations, we obtain optimized structures yielding high generation efficiencies of $2.72\times 10^{-5}$ in the pair regime, and we have detailed how both the rate at which pairs of generated photons exit the channel without loss, and the heralding efficiencies, depend on the coupling between the channel and the resonator.
@@
In future work we will extend this calculation to treat the higher levels of squeezing that are possible, and include third-order processes such as self- and cross-phase modulation that will affect the dynamics of the squeezing. In the high-squeezing regime our calculations predict significant pump depletion. As pointed out earlier [24], in this regime the entanglement between the pump and the squeezed state leads to the generation of non-Gaussian states. In future work we will extend our calculation to include the non-Gaussian dynamics between the pump, signal, and idler fields. As we have shown, high efficiency microring resonators made from III-V semiconductor materials are a promising platform for investigating non-Gaussian states generated via nonlinear interactions. Furthermore, non-Gaussian states generated this way offer a potentially more efficient approach to introducing non-Gaussianity compared with probabilistic approaches based on heralding and photon-number-resolving (PNR) detection [32].
@@ Acknowledgements
Appendix A laboratoryand ring frames
With $\chi^{(2)}(\bm{r})$ the second-order nonlinear susceptibility of the material at $\bm{r}$, take $\chi^{(2)}_{ijk}{(\bm{r})}$ and $\chi^{(2)}_{i^{\prime}j^{\prime}k^{\prime}}(\bm{r})$ to be the components of the tensor in the ring frame and the laboratoryframe respectively, with $i,j,k\in\{x,y,z\}$ and $i^{\prime},j^{\prime},k^{\prime}\in\{x^{\prime},y^{\prime},z^{\prime}\}$ (see Fig. 1). For a zincblende structure grown in the $\hat{z}$ direction, the only non-zero components of ${\chi^{(2)}}{(\bm{r})}$ for $\bm{r}\in$ ring in the laboratoryframe are:
@@
Let $\bm{M}_{\phi}$ be the matrix that transforms the coordinates from the laboratoryframe to the ring frame, by rotation of an angle $\phi$:
@@ Appendix B Degenerate SPDC
In this Appendix we treat degenerate SPDC, where both of the generated photons are within one frequency bin labelled by the centre frequency $\omega_{S}$. For variables that differ from those for nondegenerate SPDC, we will include a “hat” for the degenerate case. The Hamiltonian that describes the degenerate SPDC interaction is
@@ B.2 Pulsed excitation
B.2.1 pair regime
In the pair regime using the same approach from Sec. II.5.1, we write the generated ket (similar to [23]):
@@
where $|\hat{\beta}|^{2}/2$ is the probability of generating a pair of photons. We write the normalized two-photon ket (i.e., $\langle\hat{\mathrm{II}}\,|\,\mathopen{}\hat{\mathrm{II}}\rangle=1$)
@@
$\begin{split}\frac{|\hat{\beta}_{\lambda\lambda^{\prime}}|^{2}}{2}=&\frac{8\pi^{2}E_{P}}{(v_{P}^{\mathrm{ac}})^{2}\hbar^{3}\omega_{0}}\\ &\times\int\mathrm{d}k_{1}\mathrm{d}k_{2}\left|\hat{\mathsf{K}}_{\lambda\lambda^{\prime}}(k_{1},k_{2},X_{P})\phi\left(X_{P}\right)\right|^{2},\end{split}$ (128)
@@
The pair regime efficiency is given by
^$\begin{split}\hat{\mathcal{E}}_{\mathrm{pulse;pairregime}}^{SS}&=\frac{|\hat{\beta}|^{2}}{2|\alpha|^{2}},\end{split}$ (130)
@@
[^$\begin{split}[\hat{\mathcal{E}}_{\mathrm{pulse;pairregime}}^{SS}]_{\lambda\lambda^{\prime}}&=\frac{|\beta_{\lambda\lambda^{\prime}}|^{2}}{2|\alpha|^{2}},\end{split}$ (131)
@@
$\begin{split}\frac{|\hat{\beta}_{\lambda\lambda^{\prime}}|^{2}}{|\hat{\beta}_{\mu\mu^{\prime}}|^{2}}=\frac{[\hat{\mathcal{E}}_{\mathrm{pulse;pairregime}}^{SS}]_{\lambda\lambda^{\prime}}}{[\hat{\mathcal{E}}_{\mathrm{pulse;pairregime}}^{SS}]_{\mu\mu^{\prime}}}=\frac{\eta_{S}^{\lambda}\eta_{S}^{\lambda^{\prime}}}{\eta_{S}^{\mu}\eta_{S}^{\mu^{\prime}}}=\frac{\Gamma_{S}^{\lambda}\Gamma_{S}^{\lambda^{\prime}}}{\Gamma_{S}^{\mu}\Gamma_{S}^{\mu^{\prime}}}.\end{split}$ (132)